# Studies of stationary features in jets: 3C 279 quasar I. On-sky scattering and dynamics

### v1

@@ 4 Kinematics of quasi-stationary component
Arshakian et al. (2024) proposed a seagull-on-wave model of the QSC and showed for the BL Lac jet that the measured superluminal velocities of the QSC indicate the presence of relativistic transverse waves traveling downstream through the QSC. In this section, we examine the apparent displacements $r$ passed by the QSC during the time $\Delta t$ between consecutive observing epochs and the apparent speeds $\beta_{\rm r}=\mu D_{\rm A}(z)(1+z)/c$ expressed in units of the speed of light ($c$), where $\mu=r/\Delta t$ is the proper motion of the QSC and $D_{A}(z)$ is the angular diameter distance at redshift $z$. In fact, the apparent displacement vector of the QSC (${r}$) is a combination of the core displacement vector (${c}$) and the intrinsic displacement vector of the QSC (${s}$) (Arshakian et al. 2020). A careful analysis of the apparent displacements and observation time intervals is necessary to disentangle the statistics of intrinsic displacements of the QSC and the core.
Figure 4: The distributions of 160 apparent QSC displacements (gray area) and their uncertainties (transparent).
@@ 4.1 Apparent displacements
In VLBA maps at 15 GHz, a compact and bright radio core is used as a reference point to measure the distances to the jet features. Due to the use of the phase self-calibration procedure for reconstructing VLBI images, the absolute position information is lost. Therefore, if the core itself is moving and the QSC has its own intrinsic motion, the observer measures the apparent motion of the QSC, which is a combination of the intrinsic displacements of the core and the QSC. To study the dynamics of the QSC, we introduce the apparent displacement vector ${r}$, where the direction of the vector and its magnitude $r=\left|{\@vec{r}}\right|$ are determined by the position of the QSC between successive epochs of observations, $t$ and $t+\Delta t$, where $\Delta t$ is the time interval between successive observations. The distribution of apparent displacements has a median value of $37\pm 3.5\leavevmode\nobreak\ \mu$as, reaches a maximum at low values, and gradually decreases towards high values (Fig. 4). The errors of the apparent displacements $\delta_{\rm r}$ are estimated using the error propagation method using the uncertainties of two consecutive positions (see Arshakian et al. 2020). The median error is $\widetilde{\delta_{\rm r}}=3.6\leavevmode\nobreak\ \mu$as. We excluded the six outliers beyond 0.15 mas from further analysis. We note that an outlier with $r=0.227$ mas is measured at the longest time interval (observational gap) of 2.36 years.
The apparent displacements are measured over a wide range of observation intervals, from one day to about one year (Fig. 5), with a median observational interval of 37 days. The apparent displacements with high relative errors, $\varepsilon_{\rm r}=\delta_{\rm r}/r>0.5$, which are considered unreliable, are distributed in the lower part of the relation in the range of observational intervals from 5 to 100 days (blue color). Those with $\varepsilon_{\rm r}<0.5$ have a natural tendency to decrease as the observation interval shortens. The occurrence of large apparent displacements at high cadences ($\lesssim 10$ days) is likely attributed to strong core displacement effects.
@@
Figure 6: The selected 154 apparent displacement vectors translated to the median QSC position. The latter is marked by a plus red sign. The dashed red line is the jet central axis which connects the median position of QSC and core.
To study the orientation of apparent displacement vectors, we translated the displacement vectors so that their initial points coincide with the median center of the QSC positions (Fig. 6). Vectors of apparent QSC displacements show preferential orientation and larger magnitudes along the jet axis, very similar to the angular distribution of apparent displacements of BL Lac (Arshakian et al. 2020). We suggest that the displacement asymmetry is due to the effect of the core displacement (wobbling), which occurs preferentially along the jet axis, mainly driven by changing synchrotron opacity as new disturbances pass through the VLBI core.
To quantify displacement asymmetry, we consider azimuthal distributions of the lengths and number of displacement vectors. The azimuthal angle is the angle between the displacement vector and the downstream direction of the jet axis, which increases anticlockwise from $0\degr$ to $360\degr$. The distribution of the mean displacement lengths is calculated within an angular beam of $60\degr$ in increments $30\degr$, and the mean length errors are calculated using Eq. (1) in the Arshakian et al. (2020). The distribution of the mean lengths of displacements is significantly asymmetric, the lengths are larger along the jet ($\sim 0.6$ mas) and on average these are about three times smaller ($\sim 0.2$ mas) in direction across the jet (Fig. 7). The change in the number of vectors with azimuthal angle is similar to the behavior of the mean length (blue line). In general, the asymmetry and anisotropy of the apparent displacement vectors can be interpreted as the effect of asymmetric core displacements. The variation of the number of vectors shows significant difference in the jet direction and across to it. Moreover, the number of downstream vectors is larger by a factor of 1.5 than those upstream the jet. This is likely related to the apparent motion of the core caused by disturbances and/or inhomogeneities propagating through it. The apparent oscillatory motion of the core relative to its undisturbed position is asymmetric, typically characterized by a rapid rise in flux density (corresponding to upstream displacement vectors of the QSC as the core shifts downstream), followed by a slower decay phase (associated with downstream displacement vectors of the QSC as the core gradually returns upstream to its undisturbed position).
@@
Once we establish that the core displacements occur mainly along the jet axis, one can use the statistical method developed in Arshakian et al. (2020) to unravel the apparent displacements of QSC. It allows the evaluation of intrinsic characteristics of the core and QSC such as the mean, standard deviation, and rms (root mean square) of intrinsic displacements using projections of apparent displacements on and across the jet axis ($r_{\rm j}$ and $r_{\rm n}$, respectively; see their Eqs. (9,10), (13,14) and (17,18)). The apparent displacements that occur over small observation time intervals should display more realistic statistics of the intrinsic displacements of QSC. To examine the latter, we used the 109 apparent displacements with $\varepsilon_{\rm r}<0.5$ and considered subsamples of apparent displacements limited by observation time intervals ${\Delta t}<{\Delta t}_{\rm lim}=10,20,30,40,50,90,400$ days. For each subsample, we estimated the intrinsic characteristics of the QSC and the core (Table 1).
Table 1: Estimates of intrinsic characteristics of the QSC and radio core for cadences $\Delta t<\Delta t_{\rm lim}$ and relative errors $\varepsilon_{\rm r}<0.5$.
$\Delta t_{\rm lim}$ $N$ $\overline{s}$ $\sigma_{\rm s}$ ${\rm rms}_{\rm s}\pm\delta_{\rm rms_{\rm s}}$ $\overline{c}$ $\sigma_{\rm c}={\rm rms}_{\rm c}\pm\delta_{\rm rms_{\rm c}}$ $\overline{\Delta t}\pm\sigma_{\rm\overline{\Delta t}}$ $\overline{\beta_{\rm s}}\pm\sigma_{\rm\overline{\beta_{\rm s}}}$
@@
10 16 0.033 0.004 $0.034\pm 0.006$ $-0.012$ $0.069\pm 0.006$ $5.68\pm 2.82$ $67.84\pm 8.71$
20 30 0.028 0.013 $0.030\pm 0.007$ $-0.004$ $0.056\pm 0.002$ $10.18\pm 5.52$ $31.28\pm 4.08$
30 42 0.026 0.013 $0.029\pm 0.007$ $-0.004$ $0.050\pm 0.002$ $14.20\pm 8.10$ $21.31\pm 2.47$
40 50 0.025 0.012 $0.028\pm 0.007$ $-0.003$ $0.053\pm 0.001$ $17.85\pm 11.28$ $16.30\pm 1.83$
50 57 0.025 0.013 $0.028\pm 0.010$   $0.002$ $0.055\pm 0.002$ $21.00\pm 13.60$ $13.83\pm 1.53$
90 82 0.029 0.017 $0.034\pm 0.009$   $0.005$ $0.056\pm 0.002$ $35.28\pm 25.30$ $9.55\pm 0.97$
400 109 0.033 0.018 $0.038\pm 0.008$ $-0.001$ $0.057\pm 0.001$ $67.04\pm 70.07$ $5.76\pm 0.65$
222 Columns are as follows: (1) $\Delta t_{\rm lim}$ is the limiting value of observation time intervals (or cadences), (2) number of apparent displacements, (3) $\overline{s}$ is the mean intrinsic apparent displacement of QSC, (4) $\sigma_{\rm s}$ is the standard deviation of the intrinsic displacements, (5) ${\rm rms}_{\rm s}$ of the intrinsic displacements of QSC and its standard error, (6) $\overline{c}$ is the mean displacement of the core, (7) ${\rm rms}_{\rm c}$ of the displacements of the core and their standard errors, (8) mean and standard deviation of the cadences, (9) mean intrinsic apparent speed of QSC $\overline{\beta_{\rm s}}=\overline{s}/\overline{\Delta t}$ and its standard deviation.
If the motion of the QSC is quasi-linear, we would expect the mean intrinsic apparent displacement, $\overline{s}$, to decrease with shorter time intervals. Indeed, $\overline{s}$ decreases to a minimum value of 0.025 mas as the time interval range shortens to $\Delta t_{\rm lim}=40$ days, and then increases to 0.034 mas for even shorter intervals (Table1). We suggest that the QSC motion is directional with subsequent reversals on time scales longer than 40 days, while it becomes non-linear on characteristic time scales $\lesssim 40$ days. The latter is a subject for a separate study; here, we adopt $\Delta t_{\rm lim}=40$ days, for which the intrinsic apparent displacement reaches its minimum value of $\overline{s}=0.025$ mas. The rms displacement of the core is approximately twice that of the QSC, indicating that the anisotropic motion of the core dominates over the intrinsic motion of the QSC. This estimate remains largely unchanged over longer time intervals, fluctuating within the range of 1.6 to 1.9.
We attempted to analyse the variation of the rms of the core and QSC with epoch. However, after filtering displacements for small $\Delta t_{\rm lim}$, the statistical samples per epoch bin became insufficient to draw a definitive conclusion.
@@ 4.2 Superluminal speeds of QSC
Approximately 96% of estimated apparent speeds $\beta_{\rm r}$ of QSC exceed the speed of light and reach very high speeds $\sim 500c$ for small observation time intervals of less than 5 days (Fig. 8). The high apparent superluminal speeds are not intrinsic to the QSC, but are likely caused by core displacements occurring on timescales of $\lesssim 5$ days (black dots), as well as by large uncertainties in the apparent displacements ($\varepsilon_{\rm r}>0.5$, red dots). The apparent speed tends to decrease with increasing observational time interval. This trend is due to the QSC motion limited to the area of size determined by the scattering of QSC positions, $r_{\rm max}\sim 0.4$ mas. To show it, we plot the blue line, which represents the linear equation $\beta_{\rm r}=r_{\rm max}/(c\,\Delta t)$. The slopes of the simulated blue line and the observational points coincide well.
Regarding the intrinsic apparent speeds of the QSC, we can determine the mean intrinsic speed as $\overline{\beta}_{\rm s}=\overline{s}/\overline{\Delta t}$, where $\overline{s}$ represents the mean intrinsic displacement of the QSC, and $\overline{\Delta t}$ denotes the mean observation time interval. The smaller the observation time interval, the more accurate the estimate of the intrinsic apparent displacement of the QSC. We use the values of $\overline{s}=0.025$ mas and $\overline{\Delta t}=17.85$ days, obtained for the range of $\Delta t<\Delta t_{\rm lim}=40$ days, to estimate the mean intrinsic apparent speed of the QSC, $\overline{\beta}_{\rm s}\approx 16.3\pm 1.8$ (see Table 1). The superluminal speed of the QSC can be explained by the seagull-on-wave model, in which a relativistic transverse wave propagates through the stationary QSC, displacing it in a direction transverse to the central axis of the jet (Arshakian et al. 2024). If the relativistic transverse wave moves in a direction nearly aligned with the line of sight, the subluminal speed of the QSC in the host galaxy’s rest frame may appear superluminal in the observer’s rest frame. The fact that all measured mean intrinsic apparent speeds of the QSC are superluminal (see Table 1) provides strong evidence for the presence of a relativistic transverse wave in the 3C 279 jet.
We now attempt to estimate the characteristics of the relativistic transverse wave. Arshakian et al. (2024) derived a formula (their eq. 3) that describes the relationship between the characteristics of the QSC and the transverse wave in the reference frames of the host galaxy and the observer. This formula applies to the case where the wave propagates in the sagittal plane, the plane containing both the line of sight and the central axis of the jet:
@@
[…]
@@
We adopt $\overline{\beta}_{\rm s}=15\pm 2$ and $\sigma_{\beta_{\rm s}}=13$. To estimate the apparent speeds of transverse waves in the jet of 3C 279, a detailed analysis of the jet ridge motion is required, following the approach outlined in Cohen et al. (2014, 2015). Here, we estimate the approximate range of wave speeds in 3C 279 by utilizing the relationship between the bulk speed and transverse wave speed observed in the jet of BL Lac. In highly magnetized jets, where the magnetic field energy density exceeds the plasma energy density, the Alfvén speed can approach the speed of light, allowing transverse Alfvén waves to propagate faster than the jet’s bulk flow (e.g., Cohen et al. 2015). Assuming the presence of strong magnetic fields of the jet at the QSC distance, we assume that the Lorentz factor of the wave $\Gamma_{\rm w}\gtrsim\Gamma_{\rm beam}\approx 30$. The upper limit of the $\Gamma_{\rm w}$ of 3C 279 can be adopted from measurements of transverse patterns speeds on the BL Lac jet. Cohen et al. (2015) provides a refined estimate of the Lorentz factor for the beam as $\Gamma_{\rm beam}\approx 4.5$ of BL Lac. Additionally, superluminal speeds for Alfvén waves range between apparent velocities $\beta_{\rm{app}}\approx 3.9$ and $\beta_{\rm{app}}\approx 13.5$, with corresponding galaxy-frame speeds approximating $\beta^{g}_{\rm wave}\approx 0.98-0.998$ (or $\Gamma_{\rm w}\approx 5-15$), thus indicating a pattern speed in BL Lac to be approximately equal two or three times faster than the beam speed, $\Gamma_{\rm w}\approx\Gamma_{\rm beam}-3\Gamma_{\rm beam}$. Adopting the later approximation for the 3C 279 we obtain a range of the speeds of transverse patterns that is $\Gamma_{\rm w}\approx 30-90$, where the mean speed is $\overline{\Gamma}_{\rm w}\approx 60$ and standard deviation of the speeds $\sigma_{\Gamma_{\rm w}}\approx 10$.
@@
In Fig. 9 we plot the relation between $\beta_{\rm s}$ and $\Gamma_{\rm w}$ (Eq. 1) for a fixed $\theta=1\degr$ and different wave inclination angles ranging from $\varphi=0.03\degr-0.8\degr$ (red lines). The width of the rectangle (cyan color) represents the range $\sigma_{\Gamma_{\rm w}}=\pm 10$ from $\overline{\Gamma}_{\rm w}\approx 60$ and the length is the range $\sigma_{\beta_{\rm s}}=\pm 13$ with respect to the mean intrinsic apparent speed of QSC, $\overline{\beta}_{\rm s}\approx 15$. From these statistics we estimated the mean inclination angle of the wave $\overline{\varphi}\approx 0.25\degr$ (Eq. 1) and defined the $\varphi$ limits confined by the rectangle, $\approx 0.03\degr$ and $\approx 0.6\degr$ (see Fig. 9), which represent ±$\pm 1\sigma$ deviations from the mean inclination angle. This gives $\sigma_{\rm\varphi}\approx 0.35\degr$ and the upper limit of $\varphi\lesssim 1.3\degr$ estimated at the level $3\sigma$. It is important to note that these are rough, model-dependent estimates that may vary with changes in the jet viewing angle, the intrinsic speed of the QSC, and the properties of the transverse wave. For example, Homan et al. (2021) report a very high Doppler boosting factor of 140 and a Lorentz factor of 70 for the jet of 3C 279. If accurate, this would imply an average wave Lorentz factor of $\overline{\Gamma}_{\rm w}\approx 120$, which, in combination with $\overline{\beta}_{\rm s}\approx 15$, places an upper limit on the wave inclination angle, $\varphi\lesssim 0.5\degr$.
Figure 8: Apparent speed of the QSC as a function of the time interval between successive observation epochs. The blue line represents the lower bound for the apparent QSC speeds. Apparent speeds with high relative errors $\varepsilon_{\rm r}>0.5$ are indicated in red.
@@
Figure 9: Apparent transverse velocity of the QSC as a function of the Lorentz factor of the wave in case of the jet viewing angle $\theta=1\degr$. Different red lines correspond to the wave inclination angles $\varphi=0.03\degr,0.1\degr,0.2\degr,0.3\degr,0.4\degr,0.6\degr,0.8\degr$. The area bounded by the cyan rectangle represents the ±$\pm 1\sigma$ deviations of $\beta_{\rm s}$ and $\Gamma_{\rm w}$ from the mean values $\overline{\beta}_{\rm s}=15$ and $\overline{\Gamma}_{\rm w}=60$, respectively.

### v2

@@ 4 Kinematics of quasi-stationary component
arshakian24 proposed a seagull-on-wave model of the QSC and showed for the BL Lac jet that the measured superluminal velocities of the QSC indicate the presence of relativistic transverse waves travelling downstream through the QSC. In this section, we examine the apparent displacements $r$ passed by the QSC during the time, $\Delta t$, between consecutive observing epochs and the apparent speeds $\beta_{\rm r}=\mu D_{\rm A}(z)(1+z)/c$ expressed in units of the speed of light ($c$), where $\mu=r/\Delta t$ is the proper motion of the QSC and $D_{A}(z)$ is the angular diameter distance at redshift $z$. In fact, the apparent displacement vector of the QSC (r\@vec{r}) is a combination of the core displacement vector (c\@vec{c}) and the intrinsic displacement vector of the QSC (s\@vec{s}) (arshakian20). A careful analysis of the apparent displacements and observation time intervals is necessary to disentangle the statistics of intrinsic displacements of the QSC and the core.
Figure 4: Distributions of 160 apparent QSC displacements (light blue area) and their uncertainties (transparent).
@@ 4.1 Apparent displacements
In VLBA maps at 15 GHz, a compact and bright radio core is used as a reference point to measure the distances to the jet features. Due to the use of the phase self-calibration procedure for reconstructing VLBI images, the absolute position information is lost. Therefore, if the core itself is moving and the QSC has its own intrinsic motion, the observer measures the apparent motion of the QSC, which is a combination of the intrinsic displacements of the core and the QSC. At centimetre wavelengths, the core corresponds either to the section of the jet where the optical depth is close to unity or to a stationary feature(s) (or a combination of both) (marscher08ASPC). For a synchrotron self-absorbed core, displacements are expected to occur along the jet axis, driven by synchrotron opacity effects, as well as variations in particle density and magnetic field strength. Stationary features located near the core may appear shifted from the jet axis owing to the small viewing angle. When blended with the self-absorbed core, this can result in apparent core displacements with a transverse component relative to the jet direction.
[…]
The apparent displacements are measured over a wide range of observation intervals, from one day to about one year (Fig. 5), with a median observational interval of 37 days. The apparent displacements with high relative errors, $\varepsilon_{\rm r}=\delta_{\rm r}/r>0.5$, which are considered unreliable, are distributed in the lower part of the relation in the range of observational intervals from 5 to 100 days (blue colour). Those with $\varepsilon_{\rm r}<0.5$ have a natural tendency to decrease as the observation interval shortens. The typical positional uncertainty in Fig. 4 is 0.01 mas, meaning that we should expect displacements between two epochs to be typically of order 0.014 mas with values tailing up to three times this due to uncertainty alone. The vast majority of short cadence displacements are there likely due to simply to uncertainty, although a few of the larger displacements may be real. We caution however that large displacements over very short time intervals are suspect due both to uncertainty in the displacements themselves and in the time intervals which are accurate at best to 0.5 days.
@@
Figure 6: Selected 154 apparent displacement vectors translated to the median QSC position. The latter is marked by a plus red sign. The dashed red line is the jet central axis which connects the median position of QSC and core.
To study the orientation of apparent displacement vectors, we translated the displacement vectors so that their initial points coincide with the median centre of the QSC positions (Fig. 6). Vectors of apparent QSC displacements show preferential orientation and larger magnitudes along the jet axis, very similar to the angular distribution of apparent displacements of BL Lac (arshakian20). We suggest that the displacement asymmetry is predominantly caused by core wobbling, which occurs along the jet axis. This effect is mainly driven by variations in synchrotron opacity as new disturbances pass through the VLBI core and clearly dominates over any possible transverse displacements of the core.
To quantify displacement asymmetry, we considered azimuthal distributions of the lengths and number of displacement vectors. The azimuthal angle is the angle between the displacement vector and the downstream direction of the jet axis, which increases anticlockwise from $0\degr$ to $360\degr$. The distribution of the mean displacement lengths was calculated within an angular beam of $60\degr$ in increments $30\degr$, and the mean length errors were calculated using Eq. (1) in the arshakian20. The distribution of the mean lengths of displacements is significantly asymmetric, the lengths are larger along the jet ($\sim 0.6$ mas), and on average these are about three times smaller ($\sim 0.2$ mas) in direction across the jet (Fig. 7). The change in the number of vectors with azimuthal angle is similar to the behaviour of the mean length (blue line). In general, the asymmetry and anisotropy of the apparent displacement vectors can be interpreted as the effect of asymmetric core displacements. The variation in the number of vectors shows a significant difference in the jet direction and across to it. Moreover, the number of downstream vectors is larger by a factor of 1.5 than those upstream the jet. This is likely related to the apparent motion of the core caused by disturbances and/or inhomogeneities propagating through it. The apparent oscillatory motion of the core relative to its undisturbed position is asymmetric, typically characterized by a rapid rise in flux density (corresponding to upstream displacement vectors of the QSC as the core shifts downstream), followed by a slower decay phase (associated with downstream displacement vectors of the QSC as the core gradually returns upstream to its undisturbed position).
@@
Once we establish that the core displacements occur mainly along the jet axis, one can use the statistical method developed in arshakian20 to unravel the apparent displacements of QSC. It allows for the evaluation of intrinsic characteristics of the core and QSC such as the mean, standard deviation, and rms (root mean square) of intrinsic displacements using projections of apparent displacements on and across the jet axis ($r_{\rm j}$ and $r_{\rm n}$, respectively; see their Eqs. (9,10), (13,14), and (17,18)). The apparent displacements that occur over small observation time intervals should display more realistic statistics of the intrinsic displacements of QSC. To examine the latter, we used the 109 apparent displacements with $\varepsilon_{\rm r}<0.5$ and considered subsamples of apparent displacements constrained to cadence ranges of 0–14, 14–33, 33–60, 60–100, and 100–365 days, ensuring that the number of apparent displacements remained close to 20 (see Table 1). For the interval of 14–33 days, the intrinsic apparent displacement reaches its minimum ($\overline{s}\approx 0.02$ mas) and the corresponding ratio $\rm rms_{c}/\rm rms_{s}=1.62$.
Table 1: Estimates of intrinsic characteristics of the QSC and radio core for specified cadence ranges $\Delta T$, with relative errors $\varepsilon_{\rm r}<0.5$. [^2] Columns are as follows: (1) $\Delta T$ is the range of observation time intervals (or cadences), (2) number of apparent displacements, (3) $\overline{s}$ is the mean intrinsic apparent displacement of QSC, (4) $\sigma_{\rm s}$ is the standard deviation of the intrinsic displacements, (5) ${\rm rms}_{\rm s}$ of the intrinsic displacements of QSC and its standard error, (6) $\overline{c}$ is the mean displacement of the core, (7) ${\rm rms}_{\rm c}$ of the displacements of the core and their standard errors, (8) mean and standard deviation of the cadences, (9) mean intrinsic apparent speed of QSC $\overline{\beta_{\rm s}}=\overline{s}/\overline{\Delta t}$ and its standard deviation.
$\Delta T$ NN $\overline{s}$ $\sigma_{\rm s}$ ${\rm rms}_{\rm s}\pm\delta_{\rm rms_{\rm s}}$ $\overline{c}$ $\sigma_{\rm c}={\rm rms}_{\rm c}\pm\delta_{\rm rms_{\rm c}}$ $\overline{\Delta t}\pm\sigma_{\rm\overline{\Delta t}}$ $\overline{\beta_{\rm s}}\pm\sigma_{\rm\overline{\beta_{\rm s}}}$
@@
0 - 14 21 0.034 0.010 $035\pm 0.007 0.0$04-0.004 $0.063\pm 0.002$ $7.41\pm 4.05$ $52.42\pm 7.16$
14 - 33 22 0.019 0.010 $021\pm 0.008 0.0$04-0.004 $0.034\pm 0.004$ $21.49\pm 5.10$ $10.00\pm 1.26$
33 - 60 23 0.026 0.016 $030\pm 0.014 0.0$170.017 $0.060\pm 0.010$ $46.79\pm 8.49$ $6.38\pm 0.86$
60 - 100 22 0.038 0.019 $043\pm 0.008 0.0$05-0.005 $0.066\pm 0.004$ $80.38\pm 13.38$ $5.48\pm 0.62$
100 - 365 21 0.052 0.015 $054\pm 0.007 0.0$13-0.013 $0.058\pm 0.006$ $182.61\pm 74.12$ $3.29\pm 0.36$
If the motion of the QSC is quasi-linear, we would expect the mean intrinsic apparent displacement, $\overline{s}$, to decrease with shorter time intervals. Indeed, $\overline{s}$ decreases to a minimum value of $\approx 0.02$ mas as the time interval decreases to $\Delta T=14-33$ days, and then increases to 0.034 mas for even shorter intervals (Table 1). The latter is most likely due to uncertainties in the displacements themselves and the limited accuracy of time interval measurements of order half a day (Fig. 5). Another explanation is that the QSC motion becomes non-linear on characteristic timescales of $\lesssim 10$ days. These are a subject for a separate study; here, we adopt $\Delta T=14-33$ days, for which the intrinsic apparent displacement reaches its minimum value of $\overline{s}=0.019$ mas. The rms displacement of the core is 1.62 times larger than that of the QSC, indicating that the anisotropic motion of the core dominates over the intrinsic motion of the QSC. At longer time intervals, up to $\lesssim 100$ days, this ratio fluctuates within the range 1.6–2 (see Table 1).
We attempted to analyse the variation in the rms of the core and QSC with epoch. However, after filtering displacements for small time intervals, the statistical samples per epoch bin became insufficient to draw a definitive conclusion.
@@ 4.2 Superluminal speeds of QSC
Regarding the intrinsic apparent speeds of the QSC, we can determine the mean intrinsic speed as $\overline{\beta}_{\rm s}=\overline{s}/\overline{\Delta t}$, where $\overline{s}$ represents the mean intrinsic displacement of the QSC, and $\overline{\Delta t}$ denotes the mean observation time interval. The smaller the observation time interval, the more accurate the estimate of the intrinsic apparent displacement of the QSC. We used the values of $\overline{s}=0.019$ mas and $\overline{\Delta t}=21.5$ days, obtained for the range of $\Delta T=14-33$ days, to estimate the mean intrinsic apparent speed of the QSC, $\overline{\beta}_{\rm s}\approx 10\pm 1.3$ (see Table 1). The superluminal speed of the QSC can be explained by the seagull-on-wave model, in which a relativistic transverse wave propagates through the stationary QSC, displacing it in a direction transverse to the central axis of the jet (arshakian24). If the relativistic transverse wave moves in a direction nearly aligned with the line of sight, the subluminal speed of the QSC in the host galaxy’s rest frame may appear superluminal in the observer’s rest frame. The fact that all measured mean intrinsic apparent speeds of the QSC are superluminal (see Table 1) provides strong evidence of the presence of a relativistic transverse wave in the 3C 279 jet.
We now attempt to estimate the characteristics of the relativistic transverse wave. arshakian24 derived a formula (their eq. 3) that describes the relationship between the characteristics of the QSC and the transverse wave in the reference frames of the host galaxy and the observer. This formula applies to the case where the wave propagates in the sagittal plane, the plane containing both the line of sight and the central axis of the jet:
@@
where $\beta_{\rm s}$ is the intrinsic apparent speed of QSC as measured by observer, $\theta$ is the jet viewing angle, $\beta^{g}_{\rm s,tr}$ and $\beta^{g}_{\rm wave}$ are the intrinsic speed of the QSC and transverse wave propagation speed in the rest frame of the host galaxy, and $\varphi$ is the wave inclination angle at the QSC position with respect to the jet propagation direction. They argue that these equations are also valid for waves propagating out of the sagittal plane if the jet viewing angle is less than approximately $10\degr$. The jet viewing angle of 3C 279 is estimated to be $\lesssim 4\degr$. piner03 use 22 and 43 GHz VLBI data and report viewing angles ranging from $2\degr$ to $4\degr$, depending on the distance of the component from the jet core. rani18 estimate a jet viewing angle of $\leq 2.6\degr$ and Lorentz factor of the flow $\geq 22.4$, based on parsec-scale collimated jet observations and variability analysis at 43 GHz, while pushkarev17 derive a viewing angle of $2.4\degr$ using apparent jet speeds (lister16) and Doppler factors derived from 15 GHz flux density variability (hovatta09). Based on VLBA polarimetric observations and Doppler boosting, jorstad04 use 43 GHz VLBI data to estimate a very small viewing angle of $\approx 0.5\degr$. An analysis of 15 GHz data for the moving component C4, based on its observed kinematics and brightness evolution, suggests a required viewing angle of less than $1\degr$ and a Lorentz factor of $\Gamma>15$ (homan03). Moreover, later robust components clearly reach apparent speeds of nearly $30c$, implying a Lorentz factor of at least $\Gamma\gtrsim 30$ (homan15). The range of viewing angles could be a result of varying the jet direction as evident from changes of PAs by about $20\degr$ over 27 years, and that it may undergo nutation on shorter timescales (see Section 3), while the high Lorentz factors could be due to parsec-scale acceleration and/or bending the jet flow. The bulk Lorentz factor of the jet in 3C 279 is typically estimated to lie in the range $\Gamma_{\rm beam}\sim 10$2020 (lister13), with most studies converging around $\Gamma_{\rm beam}\sim 12$$15$ (jorstad04; homan03; piner03). In this study, we adopt a viewing angle of $\theta=1\degr$, as smaller angles would imply a jet opening angle wider than observed (pushkarev17). We also assume a bulk Lorentz factor of $\Gamma_{\rm beam}=30$, which corresponds to a maximum apparent speed of $25c$ for moving components at $\theta=1\degr$. Under these assumptions, the wave speed is constant and the inclination angle of the transverse wave remains unchanged over the observation time interval.
@@
We adopt $\overline{\beta}_{\rm s}\approx 10\pm 1.3$ and $\sigma_{\beta_{\rm s}}\approx 8$. To estimate the apparent speeds of transverse waves in the jet of 3C 279, a detailed analysis of the jet ridge motion is required, following the approach outlined in cohen14; cohen15. Here, we estimate the approximate range of wave speeds in 3C 279 by utilizing the relationship between the bulk speed and transverse wave speed observed in the jet of BL Lac. In highly magnetized jets, where the magnetic field energy density exceeds the plasma energy density, the Alfvén speed can approach the speed of light, allowing transverse Alfvén waves to propagate faster than the jet’s bulk flow (e.g. cohen15). Assuming the presence of strong magnetic fields of the jet at the QSC distance, we assume that the Lorentz factor of the wave $\Gamma_{\rm w}\gtrsim\Gamma_{\rm beam}\approx 30$. The upper limit of the $\Gamma_{\rm w}$ of 3C 279 can be adopted from measurements of transverse patterns speeds on the BL Lac jet. cohen15 provides a refined estimate of the Lorentz factor for the beam as $\Gamma_{\rm beam}\approx 4.5$ of BL Lac. Additionally, superluminal speeds for Alfvén waves range between apparent velocities $\beta_{\rm{app}}\approx 3.9$ and $\beta_{\rm{app}}\approx 13.5$, with corresponding galaxy-frame speeds approximating $\beta^{g}_{\rm wave}\approx 0.98-0.998$ (or $\Gamma_{\rm w}\approx 5-15$), indicating that the pattern speed in BL Lac is approximately two or three times faster than the beam speed, $\Gamma_{\rm w}\approx\Gamma_{\rm beam}-3\Gamma_{\rm beam}$. Adopting the later approximation for the 3C 279 we obtain a range of the speeds of transverse patterns that is $\Gamma_{\rm w}\approx 30-90$, where the mean speed is $\overline{\Gamma}_{\rm w}\approx 60$ and standard deviation of the speeds $\sigma_{\Gamma_{\rm w}}\approx 10$.
@@
In Fig. 8 we plot the relation between $\beta_{\rm s}$ and $\Gamma_{\rm w}$ (Eq. 1) for a fixed $\theta=1\degr$ and different wave inclination angles ranging from $\varphi\approx 0.03\degr-0.8\degr$ (black lines). The width of the rectangle (cyan colour) represents the range $\sigma_{\Gamma_{\rm w}}=\pm 10$ from $\overline{\Gamma}_{\rm w}\approx 60$ and the length is the range $\sigma_{\beta_{\rm s}}=\pm 8$ with respect to the mean intrinsic apparent speed of QSC, $\overline{\beta}_{\rm s}=10$. From these statistics we estimated the mean inclination angle of the wave $\overline{\varphi}\approx 0.2\degr$ (Eq. 1) and defined the $\varphi$ limits confined by the rectangle, $\approx 0.03\degr$ and $\approx 0.4\degr$ (see Fig. 8), which represent $\pm 1\sigma$ deviations from the mean inclination angle. This gives $\sigma_{\rm\varphi}\approx 0.2\degr$ and the upper limit of $\varphi\lesssim 0.8\degr$ estimated at the level $3\sigma$. It is important to note that these are rough, model-dependent estimates that may vary with changes in the jet viewing angle, the intrinsic speed of the QSC, and the properties of the transverse wave. For example, homan21 report a very high Doppler boosting factor of 140 and a Lorentz factor of 70 for the jet of 3C 279. If accurate, this would imply an average wave Lorentz factor of $\overline{\Gamma}_{\rm w}\approx 120$, which, in combination with $\overline{\beta}_{\rm s}\approx 10$, places an upper limit on the wave inclination angle, $\varphi\lesssim 0.4\degr$.
@@
Figure 8: Apparent transverse velocity of the QSC as a function of the Lorentz factor of the wave in case of the jet viewing angle $\theta=1\degr$. Different lines correspond to the wave inclination angles $\varphi=0.03\degr,0.1\degr,0.2\degr,0.3\degr,0.4\degr,0.6\degr,0.8\degr$. The area bounded by the cyan rectangle represents the $\pm 1\sigma$ deviations of $\beta_{\rm s}$ and $\Gamma_{\rm w}$ from the mean values $\overline{\beta}_{\rm s}=10$ and $\overline{\Gamma}_{\rm w}=60$, respectively.